# VideoCAD: A Large-Scale Video Dataset for Learning UI Interactions and 3D Reasoning from CAD Software

### v1

@@ Abstract
Computer-Aided Design (CAD) is a time-consuming and complex process, requiring precise, long-horizon user interactions with intricate 3D interfaces. While recent advances in AI-driven user interface (UI) agents show promise, most existing datasets and methods focus on short, low-complexity tasks in mobile or web applications, failing to capture the demands of professional engineering tools. In this work, we introduce VideoCAD, the first attempt at engineering UI interaction learning for precision tasks. Specifically, VideoCAD is a large-scale synthetic dataset consisting of over 41K annotated video recordings of CAD operations, generated using an automated framework for collecting high-fidelity UI action data from human-made CAD designs. Compared to existing datasets, VideoCAD offers an order of magnitude higher complexity in UI interaction learning for real-world engineering tasks, having up to a 20×20$\times$20 × longer time horizon than other datasets. We show two important downstream applications of VideoCAD: learning UI interactions from professional precision 3D CAD tools and a visual question-answering (VQA) benchmark designed to evaluate multimodal large language models’ (LLM) spatial reasoning and video understanding abilities. To learn the UI interactions, we propose VideoCADFormer - a state-of-the-art model in learning CAD interactions directly from video, which outperforms multiple behavior cloning baselines. Both VideoCADFormer and the VQA benchmark derived from VideoCAD reveal key challenges in the current state of video-based UI understanding, including the need for precise action grounding, multi-modal and spatial reasoning, and long-horizon dependencies. Dataset and code available at: https://github.com/BrandonMan123/VideoCAD
@@ 1 Introduction
Professional engineering design software, such as Computer-Aided Design (CAD) tools, presents significant challenges for human users due to their complex user interfaces, intricate workflows, and high-precision requirements piegl2005ten . Unlike typical consumer applications, where tasks are completed through simple UI interactions, CAD operations require structured, multi-step processes involving 3D spatial reasoning and parametric modeling. Becoming proficient in these tools can take years of expertise, and it is often very difficult to automate. Although AI systems have the potential to automate a large portion of such critical workflow for a significant productivity boost, current AI-driven automation methods struggle to generalize due to a lack of large-scale, annotated datasets for learning from CAD interfaces.
@@
Existing research in user interface navigation has largely focused on web and mobile applications, where task completion involves relatively short time horizons and does not require any 3D reasoning capabilities. Benchmarks such as MiniWob++ pmlr-v70-shi17a and RICO Deka:2017:Rico provide valuable insights into learning from UI interactions, but they lack the complexity of CAD environments, where tasks involve nested dependencies, precise geometric constraints, and tool-based interactions beyond simple button clicks. Furthermore, while behavior cloning from videos has shown promise in robotic manipulation and gaming environments, its application to software UI understanding remains underexplored due to the scarcity of high-quality, video-annotated UI datasets.
@@
To bridge this gap, we introduce VideoCAD, a large-scale dataset comprising over $41K$ synthetic video demonstrations of CAD modeling tasks and containing fine-grained action annotations. To generate VideoCAD, we leverage a public dataset of parametric CAD models (DeepCAD  wu2021deepcad ) made by humans using Onshape onshape (see Appendix A), a widely-used browser-based professional CAD platform. We map the construction sequences of CAD models into UI actions performed within OnShape.
@@
We demonstrate the effectiveness of our dataset, we provide two important downstream applications of VideoCAD. First, we develop a novel transformer-based framework, VideoCADFormer, that outperforms existing state-of-the-art methods in learning from such rich user interactions for CAD designs. We benchmark our model against other state-of-the-art behavior cloning baselines and show that VideoCADFormer is able to reason about long-horizon UI interaction tasks that involve 3D reasoning better than other models by as much as 20%. Second, we also introduce VideoCADQA, a synthetically generated multiple-choice VQA dataset derived from VideoCAD that includes questions that evaluate 3D reasoning and video understanding capabilities of LLMs. Our experiments highlight a critical gap in spatial reasoning of LLMs in precise engineering tasks.
@@
  • A VQA benchmark for 3D and spatiotemporal reasoning: We present a case study of extending VideoCAD into a VQA benchmark, introducing a suite of 1,200 visual questions that probe LLMs’ ability to perform fine-grained 3D reasoning and temporal understanding over CAD videos.
@@ 2 Related works
User Interface Navigation. Automating user interface navigation has been a long-standing challenge in AI research. Works such as MiniWob++ liu2018reinforcement and Android In The Wild (AITW) rawles2023androidwildlargescaledataset introduced datasets for web and mobile UI interaction, allowing agents to learn structured policies for performing simple tasks. However, these datasets focus on relatively simple tasks with short time horizons. More recent works such as WebShop yao2023webshopscalablerealworldweb have extended these capabilities to complex real-world scenarios where an agent must reason over long sequences of actions to achieve its goal.
@@
UI Agents and Behavior Cloning from Videos. Learning from demonstrations has been widely applied to UI interaction. Behavioral cloning, where an agent learns from human demonstrations, has been successfully applied in robotic manipulation chi2024diffusionpolicyvisuomotorpolicy and game environments baker2022videopretrainingvptlearning . However, its application to GUI-based tasks is still in its early stages. Recent studies, such as ActionBert he2021actionbertleveraginguseractions and AssistGUIs gao2024assistguitaskorienteddesktopgraphical , explored using large-scale datasets for predicting UI interactions from multimodal inputs. Our work extends these efforts by introducing a dataset specifically tailored for CAD interactions, where the action space is significantly more complex.
@@
CAD Datasets for Learning-Based Modeling. Parametric CAD datasets remain scarce compared to mesh or point cloud collections. The ABC dataset offers $1$ million B-rep models with precise geometry but lacks procedural data koch2019abc . DeepCAD supplements this by providing 178K models with full construction histories, enabling sequence-based generative modeling wu2021deepcad . Fusion 360 Gallery contributes 8.6K human-authored CAD programs, capturing sketch-extrude workflows and assembly hierarchies willis2021fusion360 . MFCAD and MFCAD++ focus on machining feature recognition, supplying labeled B-rep models for supervised learning cao2020graph ; colligan2022hierarchical . However, there is a lack of datasets that can capture the fine-grained UI interactions or the visual and geometric reasoning needed to model 3D CAD workflows. VideoCAD addresses this gap by providing video demonstrations of CAD construction, offering temporally grounded UI state changes and visual context to support multimodal learning of design behaviors.
@@
AI-Driven CAD Generation. Generative modeling of CAD has been an active research area in 3D vision Jones_2020 ; khan2024text2cadgeneratingsequentialcad ; willis2022joinablelearningbottomupassembly ; heidari2024geometricdeeplearningcomputeraided . DeepCAD wu2021deepcad introduced one of the first deep learning models capable of generating parametric CAD sequences, while subsequent works such as SketchGraphs seff2020sketchgraphslargescaledatasetmodeling have provided datasets for learning structured CAD designs. Recent advances in generative AI, such as transformer-based models for sketch generation para2021sketchgengeneratingconstrainedcad , conditional latent diffusion model for CAD sequence generation from images alam2024gencad and direct CAD construction jayaraman2023solidgenautoregressivemodeldirect , highlight the potential for learning CAD from visual demonstrations. VideoCAD complements these efforts by providing an extensive dataset of video demonstrations, enabling new paradigms in learning CAD construction from human-like interactions.
@@ 3 VideoCAD Dataset
Figure 1: Illustration of the VideoCAD dataset pipeline: human-authored CAD sequences are converted into UI instructions and executed via a rule-based bot to record videos. Quality filtering (DINOv2), keyframe extraction, and action alignment produce structured video-action pairs.
VideoCAD is a large-scale dataset $\mathcal{D}$ comprising $41{,}005$ synthetic videos of 3D CAD model construction, generated from human-authored designs. Each data sample consists of a multiframe video of an automated agent interacting with the OnShape CAD interface onshape , a free, browser-based CAD platform, accompanied by two levels of timestamped action annotations and a ground-truth target rendering in isometric view. The annotations consist of low-level actions and high-level actions. Low-level actions capture UI interactions such as clicking, typing, and mouse movements, with timestamps marking when each action occurred. These low-level actions correspond to the action space discussed in section 4. High-level actions align with CAD modeling operations, recording when primitives such as extrusions and loops are constructed. High-level actions correspond to the CAD primitives found in DeepCAD. More specifically, VideoCAD consists of the following, $\mathcal{D}=\{(\mathbf{X}_{i},\mathbf{I}_{i},\mathbf{a}_{i})\}_{i=1}^{N}$ where $\mathbf{X}_{i}\in\mathbb{R}^{t\times H\times W\times C}$ is the video frame, $\mathbf{I}_{i}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$ is the corresponding image of the target shape, and $\mathbf{a}_{i}\in\mathbb{R}^{t\times d}$ is the corresponding action vector that represents the UI actions taken to create the CAD with the desired shape.
@@ 3.1 Dataset Generation Pipeline
Sketch-extrude CAD. To create the VideoCAD dataset, we leverage the DeepCAD sequence representation wu2021deepcad , which models CAD construction as a sequence of sketches and extrusions. Each 2D sketch consists of one or more closed loops formed by geometric primitives—lines, arcs, and circles—parameterized by their spatial properties (e.g., start/end points, midpoints, radii). These sketches are extruded using parameters such as orientation (Euler angles $\theta,\phi,\gamma$), 3D offsets $(p_{x},p_{y},p_{z})$, scale $s$, bidirectional depths $(e_{1},e_{2})$, and Boolean operations $\beta$ (e.g., join, cut). This design framework is visualized in Figure 2.
Figure 2: An example of a step-by-step construction of a multi-extrusion CAD model using sketch and extrusion operations.
CAD uses a sequence of parametric commands to create the shape step by step from scratch. Although there are a lot of CAD operations in existing commercial platforms, the two most widely used CAD operations are sketch and extrusion wu2021deepcad . The purpose of the sketch command is to create the 2D sketch on a plane and then the extrusion command extrudes the sketch to create a 3D shape. To create complex and realistic 3D shapes, multiple sketches on different planes and multiple extrusion operations are required.
@@
From Human-Created CAD Models to UI instructions. We systematically convert these sequences of human-made CAD models into executable sequences of UI actions compatible with Onshape’s interface. To automate CAD modeling within Onshape, we developed a hybrid UI interaction framework driven by a rule-based bot. Each CAD model is built incrementally by defining sketch planes, either default or custom offset planes, followed by drawing geometric primitives (lines, arcs, circles), and performing extrusion operations with specified parameters. Further details on the translation from DeepCAD sequences to CAD UI actions are described in Appendix LABEL:appendix:cad_generation.
Rule-Based Bot Execution.
We use a hybrid rule-based bot to execute the instruction sequences within Onshape. The bot uses: Selenium for DOM-level browser automation, such as opening sketch menus, toggling extrusion modes, or navigating dialog boxes. PyAutoGUI for low-level pixel-based interactions within the sketch canvas, including drawing curves and typing parameter values. Since Onshape does not expose a public sketching API, this low-level simulation is essential.
All (x, y) interactions are normalized to the $(0,1)$ screen space to ensure scale invariance. A timeout-based retry mechanism (3 retries, 5s each) is used for all UI actions. Episodes are terminated after three consecutive failures—most frequently caused by interface lags or rendering issues. During execution, we record high-resolution video at 60 FPS and log every UI action with sub-second alignment, deploying the system across 64 cloud VMs. In total, our bot successfully reconstructed 44,292 CAD models out of 71,424 attempts (62.0%).
Screen Recording and Logging.
While the bot executes, we record high-resolution screen captures at 60 FPS using a custom Python video recorder. Simultaneously, we log every executed UI action (mouse event, keyboard press, selection change), each tagged with its precise frame index. This produces action–video alignment at sub-second granularity. The system is deployed across 64 Google Cloud VMs using ‘Xvfb’ to simulate GUI displays. Videos and logs are streamed to Google Cloud Storage. In total, over 118 days of screen recordings were generated in under one week.
Introducing Human-like Heuristics to VideoCAD
To balance realism and learnability in the presence of long-horizon 3D reasoning tasks (see Section 3.3), we introduce a few lightweight human-inspired heuristics into VideoCAD Buckley2018Heuristics . First, we add randomized delays between actions, with the bot waiting 0.2–0.5 seconds before interacting—mimicking human hesitation. Additionally, surface selections during sketching are made by randomly sampling a point on the surface, rather than always selecting the center. Finally, to emulate the need for precision, the bot performs zoom actions when small features are hard to select, replicating how humans zoom in to refine their input.
@@
Quality Control and Filtering. To ensure high-quality reconstruction, we render the final CAD model from an isometric view and compare it with the human-authored reference using DINOv2 vision embeddings. A cosine similarity threshold is applied to automatically discard inaccurate reproductions. We manually verified that this metric correlates well with geometric correctness (see Appendix C). After filtering, keyframes are extracted for each UI action based on the logged frame index, yielding a sequence of temporally aligned image–action pairs.
@@ 3.2 Dataset Composition and Statistics
Figure 3: Statistical distributions of CAD UI actions and sequence lengths. a. Action command frequencies. b. Sequence length frequencies.
Figure 4: (a) Overlaid kernel density estimates for the X- and Y-coordinate distributions of mouse movements. (b) Kernel density of numeric values entered via the “Type” action. (c) Relative frequencies of scroll directions (up vs. down). (d) Frequency of individual key presses, sorted from most to least common. (e) Histogram of the number of times the “tab” key is pressed.
We focus exclusively on multi-extrusion sequences from the DeepCAD dataset to create VideoCAD. Multi-extrusion sequences involve substantially longer action sequences (Figure 3) and multi-surface operations–making them more challenging for learning-based models. Each episode includes: 1) A rendered target image of shape $3\times 224\times 224$, in isometric view, 2) A full-resolution video at 1600×1000, recorded at 60 FPS, 3) A sequence of action tuples aligned with the video timeline. Figure 3 shows the action commands distribution and figure 4 shows the distribution of values of each parameter.

### v2

# VideoCAD: A Dataset and Model for Learning Long‑Horizon 3D CAD UI Interactions from Video

@@ Abstract
Computer-Aided Design (CAD) is a time-consuming and complex process, requiring precise, long-horizon user interactions with intricate 3D interfaces. While recent advances in AI-driven user interface (UI) agents show promise, most existing datasets and methods focus on short, low-complexity tasks in mobile or web applications, failing to capture the demands of professional engineering tools. In this work, we introduce VideoCAD, the first attempt to model UI interactions for precision engineering tasks. Specifically, VideoCAD is a large-scale synthetic dataset consisting of over 41K annotated video recordings of CAD operations, generated using an automated framework for collecting high-fidelity UI action data from human-made CAD designs. Compared to existing datasets, VideoCAD offers an order‑of‑magnitude increase in complexity for real‑world engineering UI tasks, with time horizons up to $20\times$ longer than those in other datasets. We show two important downstream applications of VideoCAD: (1) learning UI interactions from professional 3D CAD tools for precision tasks and (2) a visual question‑answering (VQA) benchmark designed to evaluate multimodal large language models (LLMs) on spatial reasoning and video understanding. To learn the UI interactions, we propose VideoCADFormer, a state-of-the-art model for learning CAD interactions directly from video, which outperforms existing behavior cloning baselines. Both VideoCADFormer and the VQA benchmark derived from VideoCAD reveal key challenges in the current state of video-based UI understanding, including the need for precise action grounding, multi-modal and spatial reasoning, and long-horizon dependencies. The dataset and code are available at: https://github.com/ghadinehme/VideoCAD
@@ 1 Introduction
Designing most physical products—from cars to airplanes—relies on professional engineering CAD software such as SolidWorks, Autodesk Inventor, and PTC Onshape. These platforms, with their hundreds of toolbars and menu options, pose significant challenges for users due to their complex interfaces, intricate workflows, and the high degree of precision required to create accurate 3D geometries piegl2005ten; mathur2020interactive. Unlike typical consumer applications, where tasks are completed through simple User Interface (UI) interactions, CAD operations require structured, multi-step processes involving 3D spatial understanding and parametric modeling. Mastery of these tools often requires years of experience, and automating their use remains a formidable challenge lee2025new. While recent AI-based approaches aim to automate parts of the CAD workflow, most are limited to text or image modalities you2024img2cad; khan2024text2cadgeneratingsequentialcad and fail to learn from the dynamic, video-based nature of CAD interfaces. The absence of large-scale, annotated datasets capturing such interactions further constrains progress in this area.
@@
Prior research on UI navigation has primarily focused on web and mobile applications, where tasks are short, low in complexity, and require no 3D reasoning. Benchmarks such as MiniWob++ pmlr-v70-shi17a and RICO Deka:2017:Rico have provided valuable insights into learning from user interactions, yet they fall short of capturing the complexity of CAD environments—where tasks involve nested dependencies, geometric constraints, and tool-based operations that go far beyond simple button clicks. While behavior cloning from video has shown promise in robotics and gaming, its application to software UI understanding remains largely unexplored due to the lack of large-scale, video-annotated datasets.
@@
To address this gap, we introduce VideoCAD, a large-scale synthetic dataset containing over 41K video demonstrations of CAD modeling tasks with fine-grained action annotations. VideoCAD is generated from a public dataset of parametric CAD models, DeepCAD wu2021deepcad, originally authored by human designers in Onshape onshape—a professional browser-based CAD platform (see Appendix A). We map the construction sequences of these CAD models into executable UI actions within Onshape, producing realistic, temporally aligned video–action pairs.
@@
To demonstrate the effectiveness of our dataset, we provide two important downstream applications of VideoCAD. First, we develop a transformer-based model, VideoCADFormer, that predicts UI interactions, given a single input image of target CAD model. We benchmark our model against state-of-the-art behavior cloning baselines and show that VideoCADFormer outperforms them by up to 20% in learning from rich user interactions for CAD design, demonstrating superior reasoning over long-horizon UI interaction tasks involving 3D understanding. Second, we also introduce VideoCADQA, a synthetically generated multiple-choice VQA dataset derived from VideoCAD that includes questions that evaluate 3D reasoning and video understanding in LLMs. Our benchmark helps uncover a critical gap in spatial reasoning of current LLMs in precise engineering tasks.
@@
  • A VQA benchmark for 3D and spatiotemporal reasoning: We present a case study that extends VideoCAD into a VQA benchmark, introducing 1,200 visual questions that probe LLMs’ fine‑grained 3D reasoning and temporal understanding over CAD videos.
@@ 2 Related works
Research on automating software understanding spans UI navigation, learning from demonstrations, and 3D CAD generation. We situate our work at the intersection of these areas—linking long-horizon UI modeling with spatially grounded CAD reasoning.
@@
User Interface Navigation. Automating user interface navigation has been a long-standing challenge in AI research. Works such as MiniWob++ liu2018reinforcement and Android In The Wild (AITW) rawles2023androidwildlargescaledataset introduced datasets for web and mobile UI interaction, allowing agents to learn structured policies for performing simple tasks. However, these datasets focus on relatively simple tasks with short time horizons. More recent works, such as WebShop yao2023webshopscalablerealworldweb, have extended these capabilities to complex real-world scenarios where an agent must reason over long sequences of actions to achieve its goal.
@@
UI Agents and Behavior Cloning from Videos. Learning from demonstrations has been widely applied to UI interaction. Behavioral cloning, where an agent learns from human demonstrations, has been successfully applied in robotic manipulation chi2024diffusionpolicyvisuomotorpolicy and game environments baker2022videopretrainingvptlearning. However, its application to GUI-based tasks is still underexplored. Recent studies, such as ActionBert he2021actionbertleveraginguseractions and AssistGUIs gao2024assistguitaskorienteddesktopgraphical, explored using large-scale datasets for predicting UI interactions from multimodal inputs. Our work extends these efforts by introducing a dataset specifically tailored for CAD interactions, where the action space is significantly more complex.
@@
CAD Datasets for Learning-Based Modeling. Parametric CAD datasets remain scarce compared to mesh or point cloud collections. The ABC dataset offers $1$ million B-rep models with precise geometry but lacks procedural data koch2019abc. DeepCAD supplements this by providing 178K models with full construction histories, enabling sequence-based generative modeling wu2021deepcad. Fusion 360 Gallery contributes 8.6K human-authored CAD programs, capturing sketch-extrude workflows and assembly hierarchies willis2021fusion360. MFCAD and MFCAD++ focus on machining feature recognition, supplying labeled B-rep models for supervised learning cao2020graph; colligan2022hierarchical. However, there is a lack of datasets capturing fine-grained UI interactions or the visual and geometric reasoning needed to model 3D CAD workflows. VideoCAD addresses this gap by providing video demonstrations of CAD construction, offering temporally grounded UI actions and visual context to support multimodal learning of design behaviors.
@@
AI-Driven CAD Generation. Generative modeling of CAD has been an active research area in 3D vision Jones_2020; khan2024text2cadgeneratingsequentialcad; willis2022joinablelearningbottomupassembly; heidari2024geometricdeeplearningcomputeraided. DeepCAD wu2021deepcad introduced one of the first deep learning models capable of generating parametric CAD sequences, while subsequent works such as SketchGraphs seff2020sketchgraphslargescaledatasetmodeling have provided datasets for learning structured CAD designs. Recent advances in generative AI, such as transformer-based models for sketch generation para2021sketchgengeneratingconstrainedcad, conditional latent diffusion model for CAD sequence generation from images alam2024gencad and direct CAD construction jayaraman2023solidgenautoregressivemodeldirect, highlight the potential for learning CAD from visual demonstrations. VideoCAD complements these efforts by providing an extensive dataset of video demonstrations, enabling new paradigms in learning CAD construction from human-like video demonstrations.
@@ 3 VideoCAD Dataset
Figure 1: Illustration of the VideoCAD dataset pipeline: human-authored CAD sequences are converted into UI instructions and executed via a rule-based automated method to record videos. Quality filtering, keyframe extraction, and action alignment produce structured video-action pairs.
VideoCAD is a large-scale dataset comprising 41,00541{,}005 synthetic videos of 3D CAD model construction, generated from human-authored designs. Each data sample consists of a multiframe video of an automated agent interacting with the Onshape CAD interface onshape, a free, browser-based CAD platform, accompanied by two levels of timestamped action annotations and a ground-truth target rendering in isometric view. The annotations consist of low-level actions and high-level actions. Low-level actions capture UI interactions such as clicking, typing, and mouse movements, with timestamps marking when each action occurred. These low-level actions correspond to the action space discussed in section 4. High-level actions align with CAD modeling operations, recording when primitives such as extrusions and loops are constructed. High-level actions correspond to the CAD primitives found in DeepCAD. More specifically, VideoCAD consists of the following, $\mathcal{D}=\{(\mathbf{X}_{i},\mathbf{I}_{i},\mathbf{a}_{i})\}_{i=1}^{N}$ where $\mathbf{X}_{i}\in\mathbb{R}^{t\times H\times W\times C}$ is the video frame, $\mathbf{I}_{i}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$ is the corresponding image of the target shape, and $\mathbf{a}_{i}\in\mathbb{R}^{t\times d}$ is the corresponding action vector that represents the UI actions taken to create the CAD with the desired shape.
@@ 3.1 Dataset Generation Pipeline
CAD Construction Sequence. To create the VideoCAD dataset, we leverage the DeepCAD sequence representation wu2021deepcad, which models CAD construction as a sequence of sketches and extrusions. Each 2D sketch consists of one or more closed loops formed by geometric primitives—lines, arcs, and circles—parameterized by their spatial properties (e.g., start/end points, midpoints, radii). These sketches are extruded using parameters such as orientation (Euler angles $\theta,\phi,\gamma$), 3D offsets $(p_{x},p_{y},p_{z})$, scale $s$, bidirectional depths $(e_{1},e_{2})$, and Boolean operations $\beta$ (e.g., join, cut). This design framework is visualized in Figure 2.
Figure 2: Example of intermediate modeling stages in VideoCAD. A sequence of snapshots illustrating the progressive construction of a CAD model through successive sketching and extrusion operations.
@@
From Human-Created CAD Models to UI Instructions. We systematically convert these sequences of human-made CAD models into executable sequences of UI actions compatible with Onshape’s interface. To automate CAD modeling within Onshape, we developed a hybrid UI interaction framework driven by a rule-based automated method. Each CAD model is built incrementally by defining sketch planes, either default or custom offset planes, followed by drawing geometric primitives (lines, arcs, circles), and performing extrusion operations with specified parameters. Further details on the translation from DeepCAD sequences to CAD UI actions are described in Appendix LABEL:appendix:cad_generation. A complementary discussion on the difference between high-level CAD sequences and low-level UI sequences can be found in Appendix LABEL:sec:temporal_causal
Rule-based UI Automated Method. We execute the UI instructions within the Onshape interface using a hybrid rule-based method. The mapping from high-level UI instructions to low-level UI actions is detailed in Appendix LABEL:appendix:onshape_ui. The method combines Selenium for Document Object Model (DOM)-level automation and PyAutoGUI for pixel-level input, enabling end-to-end sketching and extrusion without requiring Onshape’s internal API (Appendix LABEL:appendix:dataset-generation). To enhance realism and downstream learnability, we inject human-like heuristics, including randomized delays, surface point sampling, and zooming on small features—simulating natural user behavior in long-horizon CAD modeling tasks Buckley2018Heuristics. During execution, we record the full screen to capture the visual feedback of the automated construction of the CAD model, and we log both the low-level UI actions and the corresponding high-level CAD commands as structured video captions.
@@
Quality Control and Filtering. To ensure high-quality reconstruction, we render the final CAD model from an isometric view and compare it with the human-authored reference using DINOv2 vision embeddings. A cosine similarity threshold is applied to automatically discard inaccurate reproductions. We manually verified that this metric correlates with geometric correctness (see Appendix LABEL:vision-CAD). After filtering, keyframes are extracted for each UI action based on the logged frame index, yielding a sequence of temporally aligned image–action pairs.
@@ 3.2 Dataset Composition and Statistics
We focus exclusively on multi-extrusion sequences from the DeepCAD dataset to create VideoCAD. Multi-extrusion sequences involve substantially longer action sequences (Figure LABEL:fig:deepcad_sequence_length) and multi-surface operations–making them more challenging for learning-based models. Each episode includes: 1) A rendered target image of shape $3\times 224\times 224$, in isometric view, 2) A full-resolution video at 1600×1000, recorded at 60 FPS, 3) A sequence of action tuples aligned with the video timeline. Figure 3 shows the action commands distribution. Appendix LABEL:appendix:statistics provides additional plots on dataset statistics and representative CAD examples across complexity levels.
Figure 3: Statistical distributions of CAD UI actions and UI sequence lengths. a. Action command frequencies. b. UI sequence length frequencies.